Help me please....
A manufacturer of window frames knows from long experience that 5 percent of the production will have some type of minor defect that will require an adjustment. What is the probability that in a sample of 20 window frames:
a. None will need adjustment.
b. At least one will need adjustment.
c. More than two will need adjustment.
Thank You!

Answer by stanbon(48568)

A manufacturer of window frames knows from long experience that 5 percent of the production will have some type of minor defect that will require an adjustment. What is the probability that in a sample of 20 window frames:
a. None will need adjustment.
P(none faulty) = 20C0(0.05)^0(0.95)^20 = 1*1*0.3585=0.3585
----------------
b. At least one will need adjustment.
= 1 - P (none faulty)
= 1 - 0.3585
= 0.6415
--------------
c. More than two will need adjustment.
= 1 - P(0<=x<=2)
= 1- binomcdf(20,0.05,2)
= 0.07548
==============
Cheers,
Stan H.